I am not sure if I am even posting this word problem under the right category because I have no idea where to start. Here is the problem.
Television Sets. What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 30-in Tv set also has a width of 24 inches. What is its height.
I appreciate the help.
Thank you

Answer by checkley77(12844)

YOU HAVE A RIGHT TRIANGLE .
A^2+B^2=C^2
24^2+B^2=30^2
576+B^2=900
B^2=900-576
B^2=324
B=SQRT324
B=18 ANS. FOR THE HEIGHT.
PROOF:
24^2+18^2=900
576+324=900
900=900